# Mamba State-Space Models Are Lyapunov-Stable Learners

### v2

@@ Abstract
Mamba [21, 11] state-space models (SSMs) were recently shown to outperform state-of-the-art (SOTA) Transformer large language models (LLMs) across various tasks. Despite subsequent widespread adaptation, little work has focused on Mamba LLMs’ amenability for fine-tuning frameworks ubiquitously used for Transformer-based LLMs, e.g., mixed-precision fine-tuning (MPFT) and parameter-efficient fine-tuning (PEFT). For the former, it currently remains an open question whether Mamba’s recurrent dynamics are robust to small input changes, such as those encountered during MPFT. Using dynamical systems theory (in particular, Lyapunov exponents), we answer this question in the affirmative. We empirically validate this result through several experiments, showing that Mamba SSMs are significantly more stable to changes introduced by mixed-precision than comparable Transformers, even when both MPFT and PEFT are combined. For PEFT, we show how targeting specific memory buffers in Mamba’s customized CUDA kernels for low-rank adaptation regularizes SSM parameters, thus providing both parameter efficient learning and computational savings. Finally, with both MPFT and PEFT enabled, we explore the impact of instruction tuning Mamba SSMs for in-context learning (ICL) on natural language tasks. While pretrained Mamba and Mamba-2 models only achieve 38% and 82% (respectively) of the ICL improvements of comparable Transformer-based LLMs, we show that instruction tuning allows Mamba models to narrow this gap to 81% and Mamba-2 models to skyrocket over this gap to 132%.
**footnotetext: Correspondence to john.t.halloran@leidos.com.
@@ 1 Introduction
Innovating on previous state-space models (SSMs) [22, 10], Mamba [21, 11] has been recently proposed as an accurate, sub-quadratic alternative to Transformer large language models (LLMs). Upon their introduction in [21], Mamba SSMs were shown to greatly outperform comparable attention-based LLMs [4] across a large number of standard natural language benchmarks. Subsequently, pretrained Mamba models have been widely adapted across a large number of data modalities [36, 31, 43, 32], tasks [57, 53], and architectures  [1, 40, 34]. Despite such widespread adaptation and subsequent research threads [11, 40, 54], little work has been done to understand the amenability of Mamba SSMs for widely used fine-tuning frameworks, such as mixed-precision fine-tuning (MPFT) [38] and parameter-efficient fine-tuning (PEFT) [23, 25].
MPFT and PEFT are arguably two of the most widely utilized techniques for LLM alignment [50] and customization [51], and are typically combined to drastically decrease hardware demands needed to fine-tune modern LLMs [12]. However, direct application of MPFT for Mamba SSMs is made difficult due to potential sensitivities of Mamba’s state-space dynamics, a common concern for recurrent-based deep models [41]. To combat this, both [26] and [20] suggest full precision (FP32) may be required to perform stable training for Mamba models. Thus, it is currently an open question whether Mamba’s recurrent dynamics are stable in the presence of small input deviations, such as those introduced in MPFT.
To answer this question, we leverage theory from dynamical systems. Deriving and bounding the Lyapunov exponents for both Mamba and Mamba-2 models, we show that small input changes within the SSM layer of either model do not lead to exponentially deviating outputs. Empirically, we validate this theoretical result; compared to full-precision, deviations due to mixed-precision for Mamba inference are on par with those demonstrated by Transformer LLMs, while deviations due to MPFT are significantly more stable than those of comparable Transformers (Section 5). Furthermore, this trend continues when MPFT and PEFT are combined, where Mamba SSMs again produce significantly smaller deviations compared to comparable Transformer LLMs.
For PEFT, we show that by targeting the large memory buffers exploited by Mamba’s highly customized CUDA kernels, LoRA may be used for extremely efficient fine-tuning, while simultaneously regularizing the majority of Mamba’s SSM parameters via weight tying. We show that this leads to extremely efficient PEFT, resulting in up to 2.15 times faster training and 65.5% reduced memory compared to the largest evaluated Mamba model without MPFT or PEFT. Furthermore, this allows even the largest (2.8 billion parameter) Mamba LLMs to be fine-tuned on a single GPU with as little as 24GB of onboard memory.
Finally, using both MPFT and PEFT, we complement existing studies [40, 30] by exploring the ICL capabilities of instruction-tuned Mamba and Mamba-2 models on natural language tasks. In particular, the ICL capabilities of both pretrained Mamba and Mamba-2 models lag behind those of comparable Transformer models from the Pythia suite [4]; averaged across five standard natural language benchmarks and foundation model sizes, Mamba and Mamba-2 models only achieve 38% and 82%, respectively, of the performance improvements (relative to zero-shot) of Pythia models. However, after instruction-tuning, Mamba models are able to achieve as much as 81.5% of the average few-shot learning improvement (relative to zero-shot) of comparable Transformers, while Mamba-2 models push this to 132% of the ICL improvements achieved by Pythia models. We note that, similar to Transformer foundation models [55], (post) instruction tuning ICL appears as an emergent abilities for Mamba and Mamba-2 SSMs, manifesting for models of size 370 million parameters and greater, while failing to manifest for Mamba and Mamba-2 models of fewer parameter counts.
@@
We note that herein, when describing a particular foundation model or result, we use the term “Mamba model” to refer to one of the original models released in [21] and “Mamba-2 model” to refer to models released in [11]. While there are subtle architectural differences between these two SSMs, they share important similarities which allow our theoretical results to extend to both sets of models. In particular, Mamba and Mamba-2 models share the same state-space equations, support for SSM matrices, and design scheme of storing the majority of SSM parameters in a large memory buffer. Thus, we synonymously use the term MambaBlock to refer to the SSM layer of both Mamba and Mamba-2 models.
@@ 2 Background
Mixed-precision and parameter-efficient fine-tuning. In the current era of extremely large foundation models, both MPFT and PEFT have become ubiquitous tools for the rapid adaptation of Transformer-based LLMs towards specific applications. PEFT using adapters [23] allows a large pretrained model to be efficiently adapted for a particular downstream task by freezing the full model and training only a small number of extra parameters. Arguably the most widely used such PEFT method is LoRA [25], which injects trainable low-rank matrices into Transformer layers to approximate weight updates.
@@
To further decrease the computational demands necessary for LLM fine-tuning and inference, MPFT via mixed-precision (i.e., FP16 or BF16[27, 38] and quantized low-precision [12] have proven effective strategies to reduce GPU memory and runtime requirements without deleterious effects on downstream performance [12, 56]. Additionally, mixed-precision approaches have paved the way for hardware-aware optimizations within the self-attention module [9], greatly mitigating the quadratic complexity of Transformer LLMs. Together, PEFT and MPFT have created a rich ecosystem with which varying combinations of these approaches may be used to meet the computational constraints of a given training system. We note that post-fine-tuning quantization approaches [13] may be further used to decrease Transformer LLM computational demands, but such approaches are not considered in this work.
@@
State-space Models. Structured state-space sequence (S4) models [22, 14] are SSMs which leverage linear time-invariant (LTI) systems to combine the computational advantages of Transformers–i.e., highly parallelizable training–and recurrent neural networks (RNNs)–i.e., subquadratic autoregressive inference using recurrency. Within the S4 layer, an input signal is discretized and LTI parameters representing the input’s latent dynamics are learned. Owing to the S4 block’s latent dynamics being LTI, the S4 block’s output may be thus compactly represented as a single convolution between the input and an SSM convolution kernel (a matrix whose entries are products of LTI learnable parameters resulting from unrolling the state-space equations). However, despite hardware efficiency and long-dependency-modeling improvements, LTI-based S4 models remained inferior to Transformers of comparable parameter-sizes for natural language tasks, even when augmenting S4 layers with attention-layers for hybrid architectures [21].
@@
Innovating on these previous S4 approaches, Mamba utilizes time-varying parameters to model latent dynamics, thus broadening the ability to capture nuanced changes evolving in discrete-time. Without LTI dynamics, however, the input-output representation via the SSM convolution kernel is no longer applicable, thus voiding previous hardware-aware S4 optimizations [14]. To enable hardware efficiency with time-varying SSM parameters, [21] thus introduced extensively customized CUDA kernels which implement highly parallelized prefix sums to compute recurrent states. Subsequently, [11] considered the unrolled state-space equations and leveraged tensor contractions (i.e., einsum notation [45]) to efficiently calculate Mamba variables. The resulting Mamba-2 foundation models contained significantly larger latent-variable dimensions than the Mamba models of [21], while maintaining efficiency on modern GPU accelerators.
@@
In-context learning. ICL provides an adaptable alternative to fine-tuning. Rather than fine-tune the LLM directly, ICL augments a prompt with $n$ relevant examples (called shots) preceding the query of interest. Given sufficiently large models and pretraining data [7, 55], Transformer LLMs have proven adept at learning new concepts on the fly provided such few-shot prompting. However, it is worth noting that ICL inference time increases dramatically as the number of shots grows (due to self-attention’s quadratic complexity) and PEFT (when possible) is known to produce more accurate downstream learning results [7, 35].
@@ 3.1 Stable dynamics in the MambaBlock
The Mamba foundation models were pretrained in full FP32 precision. Consequently, official Mamba implementations have cautioned against fine-tuning or training in reduced precision [20, 26], with potential sensitivities of MambaBlock recurrent dynamics remaining an open question. We answer the latter using theory from dynamical systems. For Mamba’s discrete dynamic system in Equations 1 and  2, define
@@
The rate of divergence between two scalar $\varepsilon$-close inputs to a discrete dynamical system is bounded by the system’s maximal Lyapunov exponent $\lambda_{\mathtt{max}}{}$ [39]. Given $\lambda_{\mathtt{max}}{}$ and two initial values $(\bm{x}_{0},\mathbf{u}_{1})$ and $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)$, the maximum deviation between these points grows as [28, 47]:
$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\lambda_{\mathtt{max}}{})}).$
@@
The proof of Theorem 1 is available in Appendix B, where the maximal Lyapunov exponent for an arbitrary MambaBlock is first proven to be non-positive. The main result subsequently follows.
@@
The proof of Theorem 2 is available in Appendix C. Thus, by Theorems 1 and  2, the latent and output states of both Mamba and Mamba-2 models are stable to changes encountered during recurrency.
3.1.1 Consequences for automatic mixed-precision
During a forward pass, automatic mixed-precision (AMP) saves time and memory by computing forward activations in half-precision (FP16 or BF16). During a backward pass, AMP computes gradients in half-precision and up-casts to full-precision prior to updating. In contrast to full-precision fine-tuning, MPFT within the MambaBlock thus results in small differences to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ (which are passed through a Swish), $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), and the gradients calculated during training.
For a discrete dynamical system with $\lambda_{\texttt{max}}>0$, changes due to AMP compound after repeated expansion of the recurrent state, thus leading to exponential deviations between quantities calculated using mixed- versus full-precision. We note that Transformers are not recurrent, and thus not susceptible to such issues. Yet, just as differences introduced by quantization/mixed-precision produce output differences in Transformer results, differences are expected in Mamba results using different precision strategies. However, by Theorem 1, such differences do not exponentially compound over discrete-time within the MambaBlock.
3.2 Hardware-aware optimizations and PEFT
As matrices $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are time-varying, S4 optimizations via the SSM convolution kernel [10] are no longer applicable. However, by diagonality, each dimension may be computed in parallel. Furthermore, the recurrence along every dimension is a prefix sum (also called a scan), which is highly parallelizable [6]. [21] thus capitalizes on this through extensively customized CUDA kernels wherein the majority of temporal variables are carefully laid out in a large buffer of GPU memory and manipulated. Instantiated as a PyTorch linear layer’s weight matrix, this memory buffer $\mathbf{W}\in\mathbb{R}{}^{T\times 3d}$ is used to store and access the diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ for all $t\in\{1,\dots,T\}$, such that
$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$ (4)
where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on. The customized Mamba prefix scan kernel heavily relies on this memory layout to optimize the access pattern of $\mathbf{W}$ in Equations 1 and 2.
Similarly, Mamba-2 stores diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ in a large memory buffer $\mathbf{W}$. However, rather than utilizing the underlying recurrence to directly compute the hidden state and output at each time-step, [11] consider the matrix resulting from unrolling Equation 2 across all $t$. Mamba-2 models thus leverage the structure of the ensuing semiseparable matrix to calculate $\bm{x}_{1},\dots,\bm{x}_{T}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$ using tensor contractions, which are highly optimized on modern hardware accelerators.
When fine-tuning using LoRA, low-rank matrices are used to adapt the frozen weight matrices of targeted linear layers. The importance of $\mathbf{W}$ for both Mamba and Mamba-2 models makes it a primary candidate for LoRA adaptation. In such cases, selecting $\mathbf{W}$ for LoRA adaptation results in the following:

### v3

@@ Abstract
Mamba state-space models (SSMs) have recently outperformed state-of-the-art (SOTA) Transformer large language models (LLMs) in various tasks and been widely adapted. However, a major concern for stable learning in recurrent-based deep models (such as SSMs) is the sensitivy of their recurrent dynamics. Despite widespread adaptation, the sensitivity of Mamba’s recurrent dynamics under common fine-tuning methods–e.g., mixed-precision fine-tuning (MPFT) and parameter-efficient fine-tuning (PEFT)–remains unexplored. Empirically, we show that Mamba LLMs are extremely stable to changes introduced by combinations of MPFT and PEFT, in stark contrast to Transformer LLMs, which we demonstrate may drastically diverge from their respective full-precision counterparts under different combinations of MPFT and PEFT (despite the near-ubiquitous adaptation of these fine-tuning frameworks for attention-based models). The demonstrated robustness of Mamba LLMs are due to their recurrent dynamics, which we prove are guaranteed to be stable using dynamical systems theory (in particular, Lyapunov stability). We conclude by using MPFT and PEFT to novelly study Mamba LLMs’ in-context learning (ICL) abilities on natural language tasks, thus supplementing other recent work.
@@ 1 Introduction
Mamba(Gu & Dao, 2024a) state-space models (SSMs) have been shown to greatly outperform comparable attention-based LLMs (Biderman et al., 2023) across a large number of standard natural language benchmarks. Subsequently, pretrained Mamba models have been widely adapted across a large number of data modalities (Liu et al., 2024; Li & Chen, 2024; Quan & Li, 2024; Li et al., 2024), tasks (Xie et al., 2024; Wang et al., 2024a), and architectures  (Anthony et al., 2024; Park et al., 2024; Lieber et al., 2024). However, despite such widespread adaptation and subsequent research threads (Dao & Gu, 2024; Park et al., 2024; Wang et al., 2024b), no existing work has sought to understand the stability of learning Mamba SSMs in common LLM fine-tuning frameworks, such as mixed-precision fine-tuning (MPFT). For recurrent-based deep models, such as Mamba SSMs, the sensitivity of recurrent dynamics to small changes (e.g., those introduced by mixed-precision) is a primary concern for stable learning (Bengio et al., 1994; Pascanu et al., 2013).
To determine the stability of time-varying Mamba models, we leverage theory from dynamical systems. Under the framework of Lyapunov stability, we show that small input changes within the SSM layer of both Mamba and the more recent Mamba-2 (Dao & Gu, 2024) models do not lead to exponentially deviating outputs. Furthermore, we extend these theoretical results to account for deviations in input and latent Mamba states produced by combinations of MPFT and PEFT.
We empirically validate these theoretical results; for a large number of randomly generated SSM layers, we show that manually adjusting initial latent and input states produces maximum deviations in the output states which exponentially decrease over discrete time. Furthermore, by expanding previous divergence performance metrics (Dettmers et al., 2022; Dettmers & Zettlemoyer, 2023; Dettmers et al., 2024) and evaluating combinations of MPFT and PEFT, we show that fine-tuned Mamba LLMs do not substantially deviate in performance compared to full-precision full fine-tuning. This is demonstrated in stark contrast to comparable Transformer-based LLMs, which we show may drastically differ from their full-precision full fine-tuning counterparts, exhibiting large deviation spikes. Over two widely used fine-tuning datasets and two widely used natural language benchmarks, attention-based LLMs Pythia (Biderman et al., 2023) and OpenELM (Mehta et al., 2024) exhibit 9 and 4 large deviation spikes, respectively, whereas Mamba LLMs exhibit zero. Thus, Mamba LLMs are significantly more stable to performance deviations under MPFT and PEFT, whereas Transformer LLMs are susceptible to large deviation spikes.
Lastly, we use stable MPFT and PEFT to explore the ICL capabilities of instruction tuned Mamba and Mamba-2 models on natural language tasks, thus complementing recent studies which evaluated Mamba ICL on synthetic tasks (Park et al., 2024; Lee et al., 2024). While the ICL capabilities of pretrained Mamba models lag behind those of comparable Transformer models–Mamba and Mamba-2 models only achieve 38% and 82%, respectively, of the performance improvements (relative to zero-shot) of Pythia models–instruction tuned Mamba and Mamba-2 models improve to 81.5% and 132% of the ICL improvements achieved by Pythia. Thus, while pretrained Mamba models’ ICL capabilities lag behind comparable Transformers, we show that instruction-tuning using MPFT and PEFT greatly narrows this gap.
Our major contributions are summarized as follows:
  • We derive bounds on the Lyapunov exponents of both Mamba and Mamba-2 models’ SSM equations. Using these bounds, we theoretically show that small input changes within the MambaBlock do not lead to exponentially deviating outputs.
  • Building on the aforementioned bounds, we theoretically show that changes to MambaBlock latent and input states stemming from MPFT and PEFT do not lead to exponentially deviating outputs.
  • We empirically demonstrate the above theoretical results:
    • Directly introducing $\varepsilon$ changes to the inputs states of 100 random MambaBlock, we confirm that the resulting output states do not deviate from the unperturbed outputs over discrete time (the maximum deviation exponentially decreases).
    • For MPFT and PEFT, we measure the performance deviation between FP32 full fine-tuned and MPFT/LoRA Mamba models and compare to attention-based models. Across two fine-tuning datasets, two widely used natural language benchmarks, several model sizes, and a large number of MPFT/PEFT configurations, we show that training Mamba LLMs is significantly more stable than comparable Transformer LLMs.
  • We expand performance divergence measures previously used to study MPFT and PEFT for Transformer LLMs (Dettmers et al., 2022; Dettmers & Zettlemoyer, 2023; Dettmers et al., 2024). In addition to showing Mamba models are drastically more stable than comparable attention-based LLMs, the expanded divergence metric also demonstrates that Transformer LLMs are susceptible to large deviation spikes.
  • Using stable MPFT and PEFT, we complement recent studies by examining the ICL capabilities of Mamba/Mamba-2 models evaluated on natural language tasks. We show that instruction tuning allows SSMs to perform ICL competitively with comparable Transformer LLMs on natural language tasks.
@@
Terminology. When describing a particular foundation model or result, we use the term “Mamba model” to refer to one of the original models released in Gu & Dao (2024a) and “Mamba-2 model” to refer to models released in Dao & Gu (2024). Despite subtle architectural differences, these two SSMs share the same state-space equations and design scheme of storing the majority of SSM parameters in a large memory buffer. We thus synonymously use the term MambaBlock to refer to the SSM layer of both Mamba and Mamba-2 models.
@@ 2 Background
Mixed-precision and parameter-efficient fine-tuning. In the current era of extremely large foundation models, both MPFT and PEFT have become ubiquitous tools for the rapid adaptation of Transformer-based LLMs towards specific applications. PEFT using adapters (He et al., 2021) allows a large pretrained model to be efficiently adapted for a particular downstream task by freezing the full model and training only a small number of extra parameters. Arguably the most widely used such PEFT method is LoRA (Hu et al., 2021), which injects trainable low-rank matrices into Transformer layers to approximate weight updates.
@@
To further decrease the computational demands necessary for LLM fine-tuning and inference, MPFT via mixed-precision (i.e., FP16 or BF16(Kalamkar et al., 2019; Micikevicius et al., 2018) and quantized low-precision (Dettmers et al., 2024) have proven effective strategies to reduce GPU memory and runtime requirements without deleterious effects on downstream performance (Dettmers et al., 2024; Wu et al., 2020). Additionally, mixed-precision approaches have paved the way for hardware-aware optimizations within the self-attention module (Dao et al., 2022), greatly mitigating the quadratic complexity of Transformer LLMs. Together, PEFT and MPFT have created a rich ecosystem with which varying combinations of these approaches may be used to meet the computational constraints of a given training system.
@@
State-space Models. Structured state-space sequence (S4) models (Gu et al., 2022; Fu et al., 2023) are SSMs which leverage linear time-invariant (LTI) systems to combine the computational advantages of Transformers–i.e., highly parallelizable training–and recurrent neural networks (RNNs)–i.e., subquadratic autoregressive inference using recurrency. Within the S4 layer, an input signal is discretized and LTI parameters representing the input’s latent dynamics are learned. Owing to the S4 block’s latent dynamics being LTI, the S4 block’s output may be thus compactly represented as a single convolution between the input and an SSM convolution kernel (a matrix whose entries are products of LTI learnable parameters resulting from unrolling the state-space equations). However, despite hardware efficiency and long-dependency-modeling improvements, LTI-based S4 models remained inferior to Transformers of comparable parameter-sizes for natural language tasks, even when augmenting S4 layers with attention-layers for hybrid architectures (Gu & Dao, 2024a).
@@
Innovating on these previous S4 approaches, Mamba utilizes time-varying parameters to model latent dynamics, thus broadening the ability to capture nuanced changes evolving in discrete-time. Without LTI dynamics, however, the input-output representation via the SSM convolution kernel is no longer applicable, thus voiding previous hardware-aware S4 optimizations (Fu et al., 2023). To enable hardware efficiency with time-varying SSM parameters, (Gu & Dao, 2024a) thus introduced extensively customized CUDA kernels which implement highly parallelized prefix sums to compute recurrent states. Subsequently, Dao & Gu (2024) considered the unrolled state-space equations and leveraged tensor contractions (i.e., einsum notation (Rogozhnikov, 2022)) to efficiently calculate Mamba variables. The resulting Mamba-2 foundation models contained significantly larger latent-variable dimensions than the Mamba models of (Gu & Dao, 2024a), while maintaining efficiency on modern GPU accelerators.
@@
In-context learning. ICL provides an adaptable alternative to fine-tuning. Rather than fine-tune the LLM directly, ICL augments a prompt with $n$ relevant examples (called shots) preceding the query of interest. Given sufficiently large models and pretraining data (Brown et al., 2020; Wei et al., 2022), Transformer LLMs have proven adept at learning new concepts on the fly provided such few-shot prompting. However, it is worth noting that ICL inference time increases dramatically as the number of shots grows (due to self-attention’s quadratic complexity) and PEFT (when possible) is known to produce more accurate downstream learning results (Brown et al., 2020; Liu et al., 2022).
@@ 3 Mamba state-space models
Due to the time-variance of Equations 1 and  2, previous SSM stability results are no longer applicable (discussed further in Appendix E). The Mamba foundation models were pretrained in full FP32 precision and, subsequently, official implementations have cautioned against training in reduced precision (Gu & Dao, 2024b; Huggingface, 2024). Thus, how sensitive the MambaBlock’s recurrent dynamics are remains an open question. We next answer the latter using theory from dynamical systems.
@@ 3.1 Stable dynamics in the MambaBlock
For Mamba’s discrete dynamical system in Equations 1 and  2, define
@@
The rate of divergence between two scalar $\varepsilon$-close inputs to a discrete dynamical system is bounded by the system’s maximal Lyapunov exponent $\lambda_{\mathtt{max}}{}$ (Mikhaeil et al., 2022). Given $\lambda_{\mathtt{max}}{}$ and two initial values $(\bm{x}_{0},\mathbf{u}_{1})$ and $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)$, the maximum deviation between these points grows as (Laffargue et al., 2013; Sayama, 2015):
$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\lambda_{\mathtt{max}}{})}).$
@@
The proof of Theorem 1 is available in Appendix A, where the maximal Lyapunov exponent for an arbitrary MambaBlock is first proven to be non-positive. The main result subsequently follows.
@@
The proof of Theorem 2 is available in Appendix B. Thus, by Theorems 1 and  2, the latent and output states of both Mamba and Mamba-2 models are stable to changes encountered during recurrency.
3.1.1 Divergence bounds for automatic mixed-precision and PEFT
For an arbitrary MambaBlock, let $\bm{x}_{0},\dots,\bm{x}_{T}$ and $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ be the trajectory of latent and input states under full-precision fine-tuning. During a forward pass, automatic mixed-precision (AMP) saves time and memory by computing forward activations in half-precision (FP16 or BF16). During a backward pass, AMP computes gradients in half-precision and up-casts to full-precision prior to updating. MFPT thus introduces changes to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ (which are passed through a Swish) and latent states $\bm{x}_{0},\dots,\bm{x}_{T}$ through $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), up/downcasting of parameter updates, and the aforementioned changes to the input states.
Potentially in concert with MPFT, PEFT via LoRA induces changes by learning low-rank matrices $\hat{\mathbf{A}}_{t},\hat{\mathbf{B}}_{t},\hat{\mathbf{C}}_{t}$
$\hat{\bm{x}}_{t}$ $=(\bar{\mathbf{A}}_{t}+\hat{\mathbf{A}}_{t})\bm{x}_{t-1}+(\bar{\mathbf{B}}_{t}+\hat{\mathbf{B}}_{t})\mathbf{u}_{t}$
$\hat{\mathbf{y}}_{t}$ $=(\mathbf{C}_{t}+\hat{\mathbf{C}}_{t})\bm{x}_{t}$
Let $\bm{x}_{0}^{\prime},\dots,\bm{x}_{T}^{\prime}$ and $\mathbf{u}_{1}^{\prime},\dots,\mathbf{u}_{T}^{\prime}$ be the trajectory of latent and input states under MPFT and/or PEFT. We thus have the following two theorems: